# Cost Efficient Fairness Audit Under Partial Feedback

### v1

@@ A.4 Auxiliary Lemmas
Lemma 11 (Bretagnolle-Huber Inequality (Theorem 14.2 [35])).
@@
Let $q\in(0,\frac{1}{2})$ and $\varepsilon\in(0,1)$ such that $\frac{8q\varepsilon}{1-q}\leq 1$. Then, we have
@@
$\leq-q\left(8\varepsilon-\frac{(8\varepsilon)^{2}}{2}\right)-(1-q)\left(-\frac{8q\varepsilon}{1-q}-\frac{(8q\varepsilon)^{2}}{2(1-q)^{2}}\right)$ ($\log(1+x)\geq x-\frac{x^{2}}{2}$ and $\log(1-x)\geq-x-\frac{x^{2}}{2}$ for $x\geq 0$)
$=\frac{q(8\varepsilon)^{2}}{2}\left(\frac{q}{1-q}+1\right)$
$=\frac{32q\varepsilon^{2}}{1-q}$
$\leq 64q\varepsilon^{2}$ ($q\in(0,\frac{1}{2})$)
@@ B.1 Proof of Theorem 4
$\texttt{cost}\leq\tilde{O}\left(\sum_{a\in{\mathcal{A}}}\left(\frac{c_{feat}\beta_{a}}{q_{m,a}\varepsilon^{2}}+\sum_{y\in\{0,1\}}\frac{c_{lab}(q_{0\mid a}-p_{0\mid a})}{q_{y\mid a}}\right)\right).$
@@
$\leq\sqrt{\frac{(1-\varepsilon^{\prime})q_{y\mid a}}{(1+\varepsilon^{\prime})q_{1-y\mid a}}}\cdot\left(\frac{\varepsilon q_{m,a}}{10q_{M,a}}\right)^{2}$ (by Lemma 14)
@@
In Exp-Audit (Algorithm 2), we have set $R=\frac{3430\log(14\lvert{\mathcal{A}}\rvert/\delta)}{\min\{\widetilde{q}_{1\mid a},\widetilde{q}_{0\mid a}\}\varepsilon^{2}}$. Note that under event $\xi_{1}\cap\xi_{2}$, $\widetilde{q}_{y\mid a}\leq 1.1q_{y\mid a}$. Therefore, $\min\{\widetilde{q}_{1\mid a},\widetilde{q}_{0\mid a}\}\leq 1.1\min\{q_{1\mid a},q_{0\mid a}\}=1.1q_{m,a}$. Therefore, $R\geq\frac{3\log(14\lvert{\mathcal{A}}\rvert/\delta)}{q_{m,a}(1-\frac{\varepsilon}{36})\left(\frac{\varepsilon}{18.5}\right)^{2}}$, since $\varepsilon\in(0,1)$. This implies that in Eq. (7), if we substitute $\varepsilon_{o}=\frac{\varepsilon}{18.5}$, we have,
@@
The audit cost can be calculated by observing that the audit cost is incurred in two steps of the algorithm—first to obtain $\widetilde{q}_{y\mid a}$ and then to observe $R$ features. For estimating $\widetilde{q}_{y\mid a}$, the auditor pays $\tilde{O}\left(\frac{c_{feat}\beta_{a}+c_{lab}(q_{0\mid a}-p_{0\mid a})}{q_{y\mid a}\varepsilon^{\prime 2}}\right)$ (similar to proof of Theorem 2). Since $\varepsilon^{\prime}=\frac{1}{10}$ is a constant, the cost of estimating $\widetilde{q}_{y\mid a}$ is $\widetilde{O}\left(\frac{c_{feat}\beta_{a}+c_{lab}(q_{0\mid a}-p_{0\mid a})}{q_{y\mid a}}\right)$. Adding this cost for both $y=0$ and $y=1$ gives that the audit cost of this stage of the algorithm for a particular group $A=a$ is $\widetilde{O}\left(\sum_{y\in\{0,1\}}\frac{c_{feat}\beta_{a}+c_{lab}(q_{0\mid a}-p_{0\mid a})}{q_{y\mid a}}\right)$.
For observing $R$ features, the auditor incurs an audit cost of $R\cdot c_{feat}\cdot\beta_{a}$. Plugging $R=\widetilde{O}\left(\frac{1}{q_{m,a}\varepsilon^{2}}\right)$, we have that the audit cost for this phase is $\widetilde{O}\left(\frac{c_{feat}\beta_{a}}{q_{m,a}\varepsilon^{2}}\right)$.
@@
$=\sqrt{t}e^{\frac{L\varepsilon}{2\beta}}\int_{x\in{\mathcal{X}}}h(x)\exp\left(\left(\frac{\theta_{1}+\theta_{2}}{2}\right)^{\intercal}T(x)-W\left(\frac{\theta_{1}+\theta_{2}}{2}\right)-\frac{1}{4}\lVert\theta_{1}-\theta_{2}\rVert^{2}_{(\nabla^{2}W(\overline{\theta})+\nabla^{2}W(\widetilde{\theta}))}\right)\exp\left((\theta^{*}-\theta_{1})^{\intercal}T(x)\right)dx$ (Lemma 17, $\overline{\theta}\in[\theta_{1},\frac{\theta_{1}+\theta_{2}}{2}],\widetilde{\theta}\in[\theta_{2},\frac{\theta_{1}+\theta_{2}}{2}]$)
@@
$\leq\sqrt{t}\exp\left(\frac{L\varepsilon}{2\beta}-\frac{1}{4}\lVert\theta_{1}-\theta_{2}\rVert^{2}_{(\nabla^{2}W(\overline{\theta})+\nabla^{2}W(\widetilde{\theta}))}+(\theta^{*}-\theta_{1})^{\intercal}\mu\right)\cdot\mathbb{E}_{X\sim{\mathcal{E}}_{\frac{\theta_{1}+\theta_{2}}{2}}}\left[\exp\left((\theta^{*}-\theta_{1})^{\intercal}(T(X)-\mu\right)\right]$ ($\mu\coloneqq\mathbb{E}_{X\sim{\mathcal{E}}_{\frac{\theta_{1}+\theta_{2}}{2}}}[T(X)]$)
$\leq\sqrt{t}\cdot\exp\left(\frac{L\varepsilon}{2\beta}-\frac{1}{4}\lVert\theta_{1}-\theta_{2}\rVert^{2}_{2\kappa{\mathbf{I}}}+(\theta^{*}-\theta_{1})^{\intercal}\mu\right)\cdot\mathbb{E}_{X\sim{\mathcal{E}}_{\frac{\theta_{1}+\theta_{2}}{2}}}\left[\exp\left((\theta^{*}-\theta_{1})^{\intercal}(T(X)-\mu\right)\right]$ ($\nabla^{2}W(\theta)\succeq\kappa{\mathbf{I}}$ for all $\theta\in\Theta$)
$\leq\sqrt{t}\cdot\exp\left(\frac{L\varepsilon}{2\beta}-\frac{\kappa}{2}\lVert\theta_{1}-\theta_{2}\rVert^{2}+\lVert\theta^{*}-\theta_{1}\rVert\cdot\lVert\mu\rVert\right)\cdot\mathbb{E}_{X\sim{\mathcal{E}}_{\frac{\theta_{1}+\theta_{2}}{2}}}\left[\exp\left((\theta^{*}-\theta_{1})^{\intercal}(T(X)-\mu\right)\right]$ (Cauchy-Schwarz Inequality)
$\leq\sqrt{t}\cdot\exp\left(\frac{L\varepsilon}{\beta}-\frac{\kappa}{2}\lVert\theta_{1}-\theta_{2}\rVert^{2}\right)\cdot\mathbb{E}_{X\sim{\mathcal{E}}_{\frac{\theta_{1}+\theta_{2}}{2}}}\left[\exp\left((\theta^{*}-\theta_{1})^{\intercal}(T(X)-\mu\right)\right]$ ($\lVert\theta^{*}-\theta_{1}\rVert\leq\frac{\varepsilon}{2\beta}$; $\lVert\mu\rVert=\lVert\nabla W\left(\frac{\theta_{1}+\theta_{2}}{2}\right)\rVert\leq L$ since $W$ is $L$-Lipschitz)
$\leq\sqrt{t}\cdot\exp\left(\frac{L\varepsilon}{\beta}-\frac{\kappa}{2}\lVert\theta_{1}-\theta_{2}\rVert^{2}\right)\cdot\exp\left(\frac{\lVert\theta^{*}-\theta_{1}\rVert^{2}\lambda}{2}\right)$ (Lemma 18 as by assumption $\mathcal{B}\left(\frac{\theta_{1}+\theta_{2}}{2},\frac{1}{2\beta}\right)\subset\Theta$ and $\lVert\theta^{*}-\theta_{1}\rVert<\frac{1}{2\beta}$)
$\leq\sqrt{t}\cdot\exp\left(\frac{(L\vee\beta)\varepsilon}{\beta}+\frac{\lambda\varepsilon^{2}}{8\beta^{2}}-\frac{\kappa}{2}\lVert\theta_{1}-\theta_{2}\rVert^{2}\right)$ ($\lVert\theta^{*}-\theta_{1}\rVert\leq\frac{\varepsilon}{2\beta}$ and $L\leq(L\vee\beta)$)
Finally, we will use the assumption in the lemma statement that $\lVert\theta_{1}-\theta_{2}\rVert\geq\sqrt{\frac{3(16(L\vee\beta)\beta+\lambda)}{16\kappa\beta^{2}}\log\left(\frac{\gamma}{\varepsilon}\right)}$. Hence, we obtain
@@
$\leq\sqrt{t}\cdot\exp\left(\frac{(L\vee\beta)\varepsilon}{\beta}+\frac{\lambda\varepsilon^{2}}{8\beta^{2}}-\frac{\kappa}{2}\cdot\frac{3(16(L\vee\beta)\beta+\lambda)}{8\kappa\beta^{2}}\log\left(\frac{\gamma}{\varepsilon}\right)\right)$ ($\lVert\theta_{1}-\theta_{2}\rVert\geq\sqrt{\frac{3(16(L\vee\beta)\beta+\lambda)}{8\kappa\beta^{2}}\log\left(\frac{\gamma}{\varepsilon}\right)}$)
@@ B.3 Some Technical Lemmas
$=W\left(\frac{\theta_{1}+\theta_{2}}{2}\right)+\nabla W\left(\frac{\theta_{1}+\theta_{2}}{2}\right)^{\intercal}\left(\theta_{1}-\theta_{2}\right)+\frac{1}{2}\left\lVert\theta_{1}-\theta_{2}\right\rVert^{2}_{\nabla^{2}W(\overline{\theta})}$
@@
$W(\theta_{2})$ $=W\left(\frac{\theta_{1}+\theta_{2}}{2}\right)+\nabla W\left(\frac{\theta_{1}+\theta_{2}}{2}\right)^{\intercal}\left(\theta_{2}-\theta_{1}\right)+\frac{1}{2}\left\lVert\theta_{2}-\theta_{1}\right\rVert^{2}_{\nabla^{2}W(\widetilde{\theta})}$
@@ Appendix D Further Experiments
All codes for experiments are provided at https://github.com/nirjhar-das/fairness_audit_partial_feedback.
@@ D.2.1 Instance Construction
We vary $\varepsilon$ over the range of values $\{0.8,0.5,0.25,0.1,0.05,0.01,0.001\}$ and $(c_{feat},c_{lab})$ over the set $\{(0.5,0.25),(0.5,0.5),(0.5,1),(0.5,3)\}$. For every combination of $\varepsilon$ and $(c_{feat},c_{lab})$, we repeat the runs of the algorithm $5$ times. We compare the algorithms RS-Audit (Algorithm 1) and Exp-Audit (Algorithm 2). To be able to complete the experiment withing reasonable time, we capped $\tau$ for RS-Audit at $1000$ and $R$ in Exp-Audit at $20000$. To simulate past data, we first sample $200000$ points as follows: first sample $A\sim\mathrm{Ber}(0.7)$, then sample $Y\sim\mathrm{Ber}(q_{1\mid A})$ and then $X\sim{\mathcal{E}}_{\theta^{*}_{Y,A}}$. Here, $q_{1\mid 1}=0.7$ and $q_{1\mid 0}=0.4$. Then, we take an inference of the classifier on these points. The positively classified points were sent to the TruncEst subroutine (Line 4, Algorithm 2). This past data generation process is also repeated with every repetition of the algorithm.

### v2

@@ A.4 Auxiliary Lemmas
Lemma 11 (Bretagnolle-Huber Inequality (Theorem 14.2 [Lattimore_Szepesvári_2020])).
@@
Let $q\in(0,\frac{1}{3})$ and $\varepsilon\in(0,\frac{1}{8})$. Then, we have
@@
First note that $\frac{8q\varepsilon}{1-q}\leq\frac{1}{2}$.
@@
$\leq-q\left(8\varepsilon-\frac{(8\varepsilon)^{2}}{2}\right)-(1-q)\left(-\frac{8q\varepsilon}{1-q}-\frac{(8q\varepsilon)^{2}}{(1-q)^{2}}\right)$ ($\log(1+x)\geq x-\frac{x^{2}}{2}$ and $\log(1-x)\geq-x-x^{2}$ for $x\in(0,\frac{1}{2})$)
$=\frac{q(8\varepsilon)^{2}}{2}\left(\frac{2q}{1-q}+1\right)$
$\leq\frac{64q\varepsilon^{2}}{2}(1+1)$ ($q\in(0,1/3)$, hence $\frac{q}{1-q}\leq\frac{1}{2}$)
$\leq 64q\varepsilon^{2}$
@@ B.1 Proof of Theorem 4
$\texttt{cost}\leq\widetilde{O}\left(\underset{a\in{\mathcal{A}}}{\sum}\left(\frac{c_{feat}\mathbb{P}[f=0\mid A=a]}{q_{m,a}\varepsilon^{2}}+\underset{y\in\{0,1\}}{\sum}\frac{c_{lab}\mathbb{P}[f=0,Y=0\mid A=a]}{q_{y\mid a}}\right)\right).$
@@
$\leq\sqrt{\frac{(1+\varepsilon^{\prime})q_{1-y\mid a}}{(1-\varepsilon^{\prime})q_{y\mid a}}}\cdot\left(\frac{\varepsilon q_{m,a}}{10q_{M,a}}\right)^{2}$ (by Lemma 14)
@@
In Exp-Audit (Algorithm 2), we have set $R=\frac{3430\log(14\lvert{\mathcal{A}}\rvert/\delta)}{\min\{\widetilde{q}_{1\mid a},\widetilde{q}_{0\mid a}\}\varepsilon^{2}}$. Note that under event $\xi_{1}\cap\xi_{2}$, $\widetilde{q}_{y\mid a}\leq 1.1q_{y\mid a}$. Therefore, $\min\{\widetilde{q}_{1\mid a},\widetilde{q}_{0\mid a}\}\leq 1.1\min\{q_{1\mid a},q_{0\mid a}\}=1.1q_{m,a}$. Hence, $R\geq\frac{3\log(14\lvert{\mathcal{A}}\rvert/\delta)}{q_{m,a}(1-\frac{\varepsilon}{36})\left(\frac{\varepsilon}{18.5}\right)^{2}}$, since $\varepsilon\in(0,1)$. This implies that in Eq. (7), if we substitute $\varepsilon_{o}=\frac{\varepsilon}{18.5}$, we have,
@@
The audit cost can be calculated by observing that the audit cost is incurred in two steps of the algorithm—first to obtain $\widetilde{q}_{y\mid a}$ and then to observe $R$ features. For estimating $\widetilde{q}_{y\mid a}$, the auditor pays $\tilde{O}\left(\frac{c_{feat}\mathbb{P}[f=0\mid A=a]+c_{lab}\mathbb{P}[f=0,Y=0\mid A=a]}{q_{y\mid a}\varepsilon^{\prime 2}}\right)$ (similar to proof of Theorem 2). Since $\varepsilon^{\prime}=\frac{1}{10}$ is a constant, the cost of estimating $\widetilde{q}_{y\mid a}$ is $\widetilde{O}\left(\frac{c_{feat}\mathbb{P}[f=0\mid A=a]+c_{lab}\mathbb{P}[f=0,Y=0\mid A=a]}{q_{y\mid a}}\right)$. Adding this cost for both $y=0$ and $y=1$ gives that the audit cost of this stage of the algorithm for a particular group $A=a$ is $\widetilde{O}\left(\sum_{y\in\{0,1\}}\frac{c_{feat}\mathbb{P}[f=0\mid A=a]+c_{lab}\mathbb{P}[f=0,Y=0\mid A=a]}{q_{y\mid a}}\right)$.
For observing $R$ features, the auditor incurs an audit cost of $R\cdot c_{feat}\cdot\mathbb{P}[f=0\mid A=a]$. Plugging $R=\widetilde{O}\left(\frac{1}{q_{m,a}\varepsilon^{2}}\right)$, we have that the audit cost for this phase is $\widetilde{O}\left(\frac{c_{feat}\mathbb{P}[f=0\mid A=a]}{q_{m,a}\varepsilon^{2}}\right)$.
@@
$=\sqrt{t}e^{\frac{L\varepsilon}{2\beta}}\int_{x\in{\mathcal{X}}}h(x)\exp\left(\left(\frac{\theta_{1}+\theta_{2}}{2}\right)^{\intercal}T(x)-W\left(\frac{\theta_{1}+\theta_{2}}{2}\right)-\frac{1}{8}\lVert\theta_{1}-\theta_{2}\rVert^{2}_{(\nabla^{2}W(\overline{\theta})+\nabla^{2}W(\widetilde{\theta}))}\right)\exp\left((\theta^{*}-\theta_{1})^{\intercal}T(x)\right)dx$ (Lemma 17, $\overline{\theta}\in[\theta_{1},\frac{\theta_{1}+\theta_{2}}{2}],\widetilde{\theta}\in[\theta_{2},\frac{\theta_{1}+\theta_{2}}{2}]$)
@@
$\leq\sqrt{t}\exp\left(\frac{L\varepsilon}{2\beta}-\frac{1}{8}\lVert\theta_{1}-\theta_{2}\rVert^{2}_{(\nabla^{2}W(\overline{\theta})+\nabla^{2}W(\widetilde{\theta}))}+(\theta^{*}-\theta_{1})^{\intercal}\mu\right)\cdot\mathbb{E}_{X\sim{\mathcal{E}}_{\frac{\theta_{1}+\theta_{2}}{2}}}\left[\exp\left((\theta^{*}-\theta_{1})^{\intercal}(T(X)-\mu\right)\right]$ ($\mu\coloneqq\mathbb{E}_{X\sim{\mathcal{E}}_{\frac{\theta_{1}+\theta_{2}}{2}}}[T(X)]$)
$\leq\sqrt{t}\cdot\exp\left(\frac{L\varepsilon}{2\beta}-\frac{1}{8}\lVert\theta_{1}-\theta_{2}\rVert^{2}_{2\kappa{\mathbf{I}}}+(\theta^{*}-\theta_{1})^{\intercal}\mu\right)\cdot\mathbb{E}_{X\sim{\mathcal{E}}_{\frac{\theta_{1}+\theta_{2}}{2}}}\left[\exp\left((\theta^{*}-\theta_{1})^{\intercal}(T(X)-\mu\right)\right]$ ($\nabla^{2}W(\theta)\succeq\kappa{\mathbf{I}}$ for all $\theta\in\Theta$)
$\leq\sqrt{t}\cdot\exp\left(\frac{L\varepsilon}{2\beta}-\frac{\kappa}{4}\lVert\theta_{1}-\theta_{2}\rVert^{2}+\lVert\theta^{*}-\theta_{1}\rVert\cdot\lVert\mu\rVert\right)\cdot\mathbb{E}_{X\sim{\mathcal{E}}_{\frac{\theta_{1}+\theta_{2}}{2}}}\left[\exp\left((\theta^{*}-\theta_{1})^{\intercal}(T(X)-\mu\right)\right]$ (Cauchy-Schwarz Inequality)
$\leq\sqrt{t}\cdot\exp\left(\frac{L\varepsilon}{\beta}-\frac{\kappa}{4}\lVert\theta_{1}-\theta_{2}\rVert^{2}\right)\cdot\mathbb{E}_{X\sim{\mathcal{E}}_{\frac{\theta_{1}+\theta_{2}}{2}}}\left[\exp\left((\theta^{*}-\theta_{1})^{\intercal}(T(X)-\mu\right)\right]$ ($\lVert\theta^{*}-\theta_{1}\rVert\leq\frac{\varepsilon}{2\beta}$; $\lVert\mu\rVert=\lVert\nabla W\left(\frac{\theta_{1}+\theta_{2}}{2}\right)\rVert\leq L$ since $W$ is $L$-Lipschitz)
$\leq\sqrt{t}\cdot\exp\left(\frac{L\varepsilon}{\beta}-\frac{\kappa}{4}\lVert\theta_{1}-\theta_{2}\rVert^{2}\right)\cdot\exp\left(\frac{\lVert\theta^{*}-\theta_{1}\rVert^{2}\lambda}{2}\right)$ (Lemma 18 as by assumption $\mathcal{B}\left(\frac{\theta_{1}+\theta_{2}}{2},\frac{1}{2\beta}\right)\subset\Theta$ and $\lVert\theta^{*}-\theta_{1}\rVert<\frac{1}{2\beta}$)
$\leq\sqrt{t}\cdot\exp\left(\frac{(L\vee\beta)\varepsilon}{\beta}+\frac{\lambda\varepsilon^{2}}{8\beta^{2}}-\frac{\kappa}{4}\lVert\theta_{1}-\theta_{2}\rVert^{2}\right)$ ($\lVert\theta^{*}-\theta_{1}\rVert\leq\frac{\varepsilon}{2\beta}$ and $L\leq(L\vee\beta)$)
Finally, we will use the assumption in the lemma statement that $\lVert\theta_{1}-\theta_{2}\rVert\geq\sqrt{\frac{3(16(L\vee\beta)\beta+\lambda)}{8\kappa\beta^{2}}\log\left(\frac{\gamma}{\varepsilon}\right)}$. Hence, we obtain
@@
$\leq\sqrt{t}\cdot\exp\left(\frac{(L\vee\beta)\varepsilon}{\beta}+\frac{\lambda\varepsilon^{2}}{8\beta^{2}}-\frac{\kappa}{4}\cdot\frac{6(16(L\vee\beta)\beta+\lambda)}{8\kappa\beta^{2}}\log\left(\frac{\gamma}{\varepsilon}\right)\right)$ ($\lVert\theta_{1}-\theta_{2}\rVert\geq\sqrt{\frac{3(16(L\vee\beta)\beta+\lambda)}{8\kappa\beta^{2}}\log\left(\frac{\gamma}{\varepsilon}\right)}$)
@@ B.3 Some Technical Lemmas
$=W\left(\frac{\theta_{1}+\theta_{2}}{2}\right)+\nabla W\left(\frac{\theta_{1}+\theta_{2}}{2}\right)^{\intercal}\left(\theta_{1}-\theta_{2}\right)+\frac{1}{8}\left\lVert\theta_{1}-\theta_{2}\right\rVert^{2}_{\nabla^{2}W(\overline{\theta})}$
@@
$W(\theta_{2})$ $=W\left(\frac{\theta_{1}+\theta_{2}}{2}\right)+\nabla W\left(\frac{\theta_{1}+\theta_{2}}{2}\right)^{\intercal}\left(\theta_{2}-\theta_{1}\right)+\frac{1}{8}\left\lVert\theta_{2}-\theta_{1}\right\rVert^{2}_{\nabla^{2}W(\widetilde{\theta})}$
@@ Appendix D Further Experiments
All codes for experiments can be found at https://github.com/nirjhar-das/fairness_audit_partial_feedback. All codes for experiments can be found in the supplementary file.
@@ D.2.1 Instance Construction
We set $\gamma=0$ vary $\varepsilon$ over the range of values $\{0.8,0.5,0.25,0.1,0.05,0.01,0.001\}$ and $(c_{feat},c_{lab})$ over the set $\{(0.5,0.25),(0.5,0.5),(0.5,1),(0.5,3)\}$. For every combination of $\varepsilon$ and $(c_{feat},c_{lab})$, we repeat the runs of the algorithm $5$ times. We compare the algorithms RS-Audit (Algorithm 1) and Exp-Audit (Algorithm 2). To be able to complete the experiment withing reasonable time, we capped $\tau$ for RS-Audit at $1000$ and $R$ in Exp-Audit at $20000$. To simulate past data, we first sample $200000$ points as follows: first sample $A\sim\mathrm{Ber}(0.7)$, then sample $Y\sim\mathrm{Ber}(q_{1\mid A})$ and then $X\sim{\mathcal{E}}_{\theta^{*}_{Y,A}}$. Here, $q_{1\mid 1}=0.7$ and $q_{1\mid 0}=0.4$. Then, we take an inference of the classifier on these points. The positively classified points were sent to the TruncEst subroutine (Line 4, Algorithm 2). This past data generation process is also repeated with every repetition of the algorithm.
Adult Income Dataset.
With the Adult Income dataset, we first use $70\%$ of the dataset to train a 4-layer neural network with hidden layers having dimensions $64,32$ and $10$ ($64$ is closer to input layer and $10$ is closer to output layer) to predict the true labels. We train the neural net with Adam [29] otpimizer for $1000$ epochs. Thereafter, we obtain the embedding of the entire dataset by removing the last layer of the neural network and inferring on it. The resulting embedding is $10$ dimensional. This way, we map the whole dataset to the embedding space. Hereon, by the term dataset, we will refer to this embedded dataset rather than the original one. We sample random rows from the dataset to simulate online data. We first use $1000$ random samples to train a logistic regression classifier which serves as our classifier $f$ to be audited. For simulating past database, we sample $20000$ random rows from the dataset and collect a subset of those rows that correspond to $f=1$. Note that in this experiment, no structural or separation assumptions are enforced.
We set various choices of $(\gamma,\varepsilon)$, namely, $\{(0.5,0.3),(0.4,0.2),(0,0.1),(0,0.05),(0,0.01)\}$. The values of $(c_{feat},c_{lab})$ we experiment with are $\{(0.5,0.25),(0,1),(0.5,1)\}$. We repeat every choice of $((\gamma,\varepsilon),(c_{feat},c_{lab}))$ over $5$ runs and report the average and standard deviations. We compare RS-Audit and Exp-Audit. To be able to complete the experiment withing reasonable time, we capped $\tau$ for RS-Audit at $1000$ and $R$ in Exp-Audit at $20000$.